Fire fighters sprayed a 9 inch thick layer of foam over a burning rectangle region 10 yards wide and 48 feet long .what volume of foam was used on the fire ?
A 40 cu yd B 120 cu yd C 240 cu yd D 480 cu yd E 4500 cu yd

Answer by mananth(16946)

9"= 9/ 12*3= 1/4 yard
the length = 48 feet = 16yards
width =10 yards
..
Volume of foam = 1/4 * 16 * 10
= 40 cu.yard